POLYNOMIAL RETRACTS AND THE JACOBIAN CONJECTURE

VLADIMIR SHPILRAIN JIE-TAI YU

Abstract. Let K[x, y] be the polynomial algebra in two variables over a field K of characteristic 0. A subalgebra R of K[x, y] is called a retract if there is an idempotent homomorphism (a retraction, or projection) φ : K[x, y] → K[x, y] such that φ(K[x, y]) = R. The presence of other, equivalent, definitions of retracts provides several different methods of studying them, and brings together ideas from combinatorial algebra, homological algebra, and algebraic geometry. In this paper, we characterize all the retracts of K[x, y] up to an automorphism, and give several applications of this characterization, in particular, to the well-known Jacobian conjecture. Notably, we prove that if a polynomial mapping φ of K[x, y] has invertible Jacobian matrix and fixes a non-constant polynomial, then φ is an automorphism.